# Rethinking 3D Dense Caption and Visual Grounding in A Unified Framework through Prompt-based Localization

### v1

@@ 4.5 Qualitative Comparison
As for the 3DVG task, we compare qualitative results to the state-of-the-art model EDA (Wu et al., 2023) in the left part of Figure 3. Our method produces the right bounding boxes in some difficult samples, which indicates that by jointly training the 3DDC and 3DVG tasks, the 3DVG task benefits from the 3DDC task’s comprehensive characterization of object attributes. As for the 3DDC task, we compare qualitative results to the state-of-the-art model Vote2Cap-DETR (Chen et al., 2023a) in the right part of Figure 3. Our method can produce accurate descriptions of object attributes, classes, and spatial relationships. This suggests the 3DDC task benefits from the 3DVG task’s grasp of the relationships between objects in the scene to generate more accurate caption information. Through qualitative comparison, we can more intuitively recognize the advantages of an end-to-end training scheme compared to the two-stage approach.
5 Conclusions
In conclusion, we introduce a significant shift in 3DVG and 3DDC integration using the proposed 3DGCTR. By rethinking prompt-based localization in 3DVG to enhance 3DDC, this approach moves away from traditional two-stage, detector-dependent methods towards a more effective, unified strategy. The DETR-like framework enables end-to-end training, utilizing shared modules for multitasking. 3DGCTR’s exceptional performance on ScanRefer and its ability to mutually enhance 3DVG and 3DDC tasks underline its transformative impact on 3D scene understanding.

### v2

@@ 4.5 Qualitative Comparison
As for the 3DVG task, we compare qualitative results to the state-of-the-art model EDA [34] in the left part of Figure 3. Our method produces the right bounding boxes in some difficult samples, which indicates that by jointly training the 3DDC and 3DVG tasks, the 3DVG task benefits from the 3DDC task’s comprehensive characterization of object attributes. As for the 3DDC task, we compare qualitative results to the state-of-the-art model Vote2Cap-DETR [8] in the right part of Figure 3. Our method can produce accurate descriptions of object attributes, classes, and spatial relationships. This suggests the 3DDC task benefits from the 3DVG task’s grasp of the relationships between objects in the scene to generate more accurate caption information. Through qualitative comparison, we can more intuitively recognize the advantages of an end-to-end training scheme compared to the two-stage approach.
Figure 3: Qualitative Comparisons. We compare qualitative results with two state-of-the-art methods in 3DVG (left part in the figure) and 3DDC (right part in the figure) tasks, EDA [34] and Vote2Cap-DETR [8]. We mark correct attribute words in green and wrong descriptions in red. Our method produces right bounding boxes close to ground truth annotations and produces accurate descriptions of object attributes, classes and spatial relationships.
5 Limitations & Conclusions
Limitations. The main limitation is scene sampling. Small objects with few sampled points may lead to a compromised understanding of detailed features. Designed for scenes with fixed limited sampled points, the model performs well in most indoor settings but may struggle with larger scenes. Conclusions. We introduce a significant shift in 3DVG and 3DDC integration using the proposed 3DGCTR. By rethinking prompt-localization in 3DVG to enhance 3DDC, this approach moves away from traditional two-stage, detector-dependent methods towards a more effective, unified strategy. The DETR-like framework enables end-to-end training, utilizing shared modules for multitasking. 3DGCTR’s exceptional performance on ScanRefer and its ability to mutually enhance 3DVG and 3DDC tasks underline its transformative impact on 3D scene understanding.